OpenStudy (anonymous):

Evaluate: C = 5/9 for (F-32) For F = 77 degrees

OpenStudy (anonymous):

Firstly tell : \(77 - 32 = ??\)

OpenStudy (anonymous):

that equals 45

OpenStudy (anonymous):

Yep: Now do solve this: \[C = \frac{5 \times 45}{9}\]

OpenStudy (anonymous):

25

OpenStudy (nathan917):

c = 5/9(77 - 32) c = 5/9(45) c = 25

OpenStudy (anonymous):

thank you

OpenStudy (nathan917):

Anytime.